The daughter is 36 years younger than her mother. How many years ago was the mother 5 times the daughters age, if she is 50 years old now?
Found 3 solutions by Alan3354, ankor@dixie-net.com, Edwin McCravy:
Answer by Alan3354(69443)

The daughter is 36 years younger than her mother. How many years ago was the mother 5 times the daughters age, if she is 50 years old now?
------------
The difference is constant at 36 years.
When the mother was 5x, the difference was 4x
4x = 36
x = 9
------------
Can you finish it?

Answer by ankor@dixie-net.com(22740)

The daughter is 36 years younger than her mother.
How many years ago was the mother 5 times the daughters age, if she is 50 years old now?
daughter's age: 50 - 36 = 14 yrs
let y = no. of years for this to true
50 - y = 5(14 - y)
50 - y = 70 - 5y
-y + 5y = 70 - 50
4y = 20
y = 20/4
y = 5 yrs ago
:
;
Check
50 - 5 = 5(14 - 5)
45 = 5(9)
:


Answer by Edwin McCravy(20054)

Read this part first:

if she (Mother) is 50 years old now?
M = 50

The daughter is 36 years younger than her mother.
 D = M - 36
 D = 50 - 36
 D = 14

How many years ago was the mother 5 times the daughters age,
  M - Y = 5(D - Y)
 50 - Y = 5(14 - Y)
 50 - Y = 70 - 5Y
    +5Y      + 5Y
 ----------------
 50 +4Y = 70
-50      -50
------------
     4Y = 20
      Y = 20/4
      Y = 5 years ago.

Checking: 5 years ago the Mother was 45 and the Daughter was 9
and indeed 45 is 5 times 9.

Edwin